OpenStudy (anonymous):

Determine which equation below has a solution of a = 2. A. 2a + 3a – 6 = –21 B. 8 = 3 + 4a – 3 – 5a C. –2 = 2 + 5a – 2 – 6a D. –4a + 3a – 6 = –4

OpenStudy (anonymous):

C. –2 = 2 + 5a – 2 – 6a 6a-5a=2-2+2 a=2